HIGH RESOLUTION INTRACAVITY LASER SPECTRA OF NICKEL HYDRIDE

\underline{SADASIVAN SHAJI}, JAMES J. O’BRIEN, JOSHUA NUNN, MOYANG LI, ALEX SONG, Department of Chemistry and Center for Molecular Electronics, University of Missouri, St. Louis, MO 63121-4499; LEAH C. O’BRIEN, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652.

The visible electronic spectrum of the nickel hydride (NiH) is recorded with rotational resolution by intracavity laser absorption spectroscopy (ILS). A dye laser based ILS system is used to record the spectra. The gas phase NiH molecules were produced in an electric discharge using a nickel hollow cathode in a pure hydrogen atmosphere at 2-3 torr total pressure. Transitions are observed from the v=0 X ^{2}Δ_{5/2} state to various vibrational levels of several excited states, including the A ^{2}Δ_{5/2}, B ^{2}Δ_{5/2},^{2}Δ_{3/2}, and ^{2}Φ_{7/2} states. Peak positions for the NiH isotopomers will be presented for transitions involving these levels.